OpenStudy (anonymous):

a) y=logx+100 b) y=3x-6 Find another way to rewrite each one with a different sequence of transformations

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iambatman

ganeshie8 (ganeshie8):

for a, try this : y = logx + 100 = logx + log 10^100 = log(10^100x)

ganeshie8 (ganeshie8):

you can think of it as a horizontal compression by a factor of 10^100

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

for the second one, y=3x-6, how would you create a different transformation

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

try below two transformations : 1) y = 3(x-0) - 6 2) y = 3(x-2) + 0

ganeshie8 (ganeshie8):

first transofrmation represents : vertical stretch by 3 and shift down by 6 units

ganeshie8 (ganeshie8):

second transformation represents : shift right by 2 units, vertical stretch by 3

OpenStudy (anonymous):

thanks